OpenStudy (anonymous):

will medal and fan Complete the equation for the linear function whose graph contains the point (7, 8) and has a slope of −1.2 y=___x+___

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@shrutipande9 @undeadknight26

OpenStudy (xapproachesinfinity):

we need to use point slope form y=mx+b m=-1.2 so y=-1.2x+b we want to find b now if the point (7,8) is on this line then we have: 8=-1.2(7)+b solve for b

OpenStudy (xapproachesinfinity):

can you find b?

OpenStudy (anonymous):

0.4?

OpenStudy (xapproachesinfinity):

how?

OpenStudy (xapproachesinfinity):

show all the work

OpenStudy (anonymous):

im lagging so it mite take a lil

OpenStudy (xapproachesinfinity):

no problem waiting here :)

OpenStudy (anonymous):

well its hard to explain but thats what i got y=-1.2x+0.4

OpenStudy (xapproachesinfinity):

why is it hard to explain: \[\large 8=-1.2\times7+b\] \[\large 8=-8.4+b\] \[\large b=8+8.4=16.4\] i got a different answer!!

OpenStudy (xapproachesinfinity):

its just a simple linear equation?!

OpenStudy (anonymous):

y=1.2x+-0.4

OpenStudy (xapproachesinfinity):

you subtract add and stuff like that

OpenStudy (xapproachesinfinity):

what did you do to get 0.4?

OpenStudy (anonymous):

i just graphed mine it goes through (7,8)

OpenStudy (xapproachesinfinity):

no that's not the answer the slope is given to you its -1.2=m b=16.4 as i just did

OpenStudy (anonymous):

the one you came up with does not go through (7,8)

OpenStudy (anonymous):

both of ours do

OpenStudy (anonymous):

im gonna go with yours your probably way better than me at math lol

OpenStudy (xapproachesinfinity):

why would you graph? say that your answer is correct y=-1.2x+0.4 then if plugging 7 for x i should get 8 let's see if that's true y=-1.2(7)+0.4=-8.4+0.4=-8

OpenStudy (xapproachesinfinity):

no yours is incorrect dear friend

OpenStudy (xapproachesinfinity):

(7,8) is not in your line that you did y=-1.2x+0.4

OpenStudy (xapproachesinfinity):

no i'm not just make sure what you are doing is correct how would you graph is you don't have the equation in the first place?? always follow common sense

OpenStudy (xapproachesinfinity):

if*

OpenStudy (xapproachesinfinity):

you can't graph something that you don't know its equation!!!

OpenStudy (xapproachesinfinity):

i'm lagging again darn hehe

OpenStudy (anonymous):

ok thank you for your help

OpenStudy (xapproachesinfinity):

welcome!